ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 96, 1996

Agnarsson, G., S. A. Amitsur and J. C. Robson Recognition of matrix rings II

1

Amitsur, S. A. See G. Agnarsson

Amitsur, S. A. and D. Baum Finite-dimensional subalgebras of division rings

15

Amitsur, S. A. and L. W. Small Algebras over infinite fields, revisited

23

Bahturin, Y., D. Fischman and S. Montgomery On the generalized Lie structure of associative algebras

27

Baum, D. See S. A. Amitsur

Berele, A., A. Giambruno and A. Regev Involution codimensions and trace codimensions of matrices are asymptotically equal

49

Bergman, G. M. On the growth of algebras with bialgebra action  63

Bokut’, L. and P. Malcolmson Gröbner–Shirshov bases for quantum enveloping algebras

97

Braun, A. Localization, completion and the AR property in Noetherian P.I. rings

115

Brussel, E. S. Decomposability and embeddability of discretely Henselian division algebras

141

Cohen, M., S. Westreich and S. Zhu Determinants, integrality and Noether’s theorem for quantum commutative algebras

185

Cohn, P. M. On skew field coproducts with a finite extension  223

Drensky, V. and A. Regev Exact asymptotic behaviour on the codimensions of some P.I. algebras

231

Fein, B. and M. Schacher Sums of corestrictions of cyclic algebras 243

Fischman, D. See Y. Bahturin

Ford, T. J. The Brauer group of a curve over a strictly local discrete valuation ring

259

Frumkin, A. The irreducible characters of the Lie superalgebras of type A(n, m) and filtrations of their Kac modules

267

Giambruno, A. and A. Valenti Central polynomials and matrix invariants

281

Giambruno, A. See A. Berele Haile, D. E., M-A. Knus, M. Rost and J-P. Tignol Algebras of

odd degree with involution, trace forms and dihedral extensions

299

Keller, J. See V. K. Kharchenko

Kemer, A. Remarks on the prime varieties  341

Kharchenko, V. K., J. Keller and S. Rodrígues-Romo Prime rings with PI rings of constants

357

Knus, M-A. See D. E. Haile Lam, T. Y. and A. Leroy Recognition and computations of matrix

rings

379

Leroy, A. See T. Y. Lam Lubotzky, A., L. Pyber and A. Shalev Discrete groups of slow

subgroup growth

399

Makar-Limanov, L. On the hypersurface x + x^{2}y + z^{2} + t^{3} = 0 in C^{4} or a C^{3}-like threefold which is not C^{3}

419

Malcolmson, P. See L. Bokut’ Malle, G. and J. Sonn Covering groups of almost simple groups as

Galois groups over ℚ^{ab}(t)

431

Mann, A. Shimshon Avraham Amitsur (1921–1994)  ix
Mann, A. Philip Hall’s ‘rather curious’ formula for abelian p-groups 445

Mann, A. and A. Shalev Simple groups, maximal subgroups, and probabilistic aspects of profinite groups

449

Martinez, C. and E. Zelmanov Products of powers in finite simple groups

469

Montgomery, S. See Y. Bahturin Passman, D. S. The semiprimitivity problem for group algebras of

locally finite groups

481

Plotkin, B. Varieties of algebras and algebraic varieties  511

Pyber, L. See A. Lubotzky Regev, A. See A. Berele Regev, A. See V. Drensky

E. Rips On a double of a free group  523

Robson, J. C. See G. Agnarsson Rodrígues-Romo, S. See V. K. Kharchenko Rost, M. See D. E. Haile Rowen, L. H. and D. J. Saltman Semidirect product division

algebras

527

Rowen, L. H. and J-P. Tignol On the decomposition of cyclic algebras

553

Saltman, D. J. See L. H. Rowen Schacher, M. See B. Fein Shalev, A. See A. Lubotzky Shalev, A. See A. Mann Small, L. W. See S. A. Amitsur Sonn, J. See G. Malle Tignol, J-P. See D. E. Haile Tignol, J-P. See L. H. Rowen Valenti, A. See A. Giambruno Van Oystaeyen, F. and Y. H. Zhang The crossed coproduct

theorem and Galois cohomology

579

Westreich, S. See M. Cohen

Zalesski, A. E. Modular group rings of the finitary symmetric group 609

Zelmanov, E. See C. Martinez Zhang, Y. H. See F. Van Oystaeyen Zhu, S. See M. Cohen